OpenStudy (anonymous):

Find the Cartesian equation of the curve given by the parametric equations: x=(1-t^2)/(1+t^2) and y=2t/(1+t^2). I know how to do Cartesian equations, but I am unsure how to solve for t in this one.

OpenStudy (anonymous):

\[\begin{align*}\cos y&=\cos\left(2\times\frac{y}{2}\right)\\\\ &=\cos^2\frac{y}{2}-\sin^2\frac{y}{2}\\\\ \sin y&=\sin\left(2\times\frac{y}{2}\right)\\\\ &=2\sin\frac{y}{2}\cos\frac{y}{2}\end{align*}\] Recall that \(\cos^2y+\sin^2y=1\), so we should have \[1=\left(\cos^2\frac{y}{2}-\sin^2\frac{y}{2}\right)^2+\left(2\sin\frac{y}{2}\cos\frac{y}{2}\right)^2\] Suppose we have \[\cos\frac{y}{2}=\frac{1}{\sqrt{1+t^2}}~~~~\text{and}~~~~\sin\frac{y}{2}=\frac{t}{\sqrt{1+t^2}}\] Then indeed, we have \[\begin{align*}1&=\left(\cos^2\frac{y}{2}-\sin^2\frac{y}{2}\right)^2+\left(2\sin\frac{y}{2}\cos\frac{y}{2}\right)^2\\\\ &=\left(\left(\frac{1}{\sqrt{1+t^2}}\right)^2-\left(\frac{t}{\sqrt{1+t^2}}\right)^2\right)^2+\left(2\times\frac{1}{\sqrt{1+t^2}}\times\frac{t}{\sqrt{1+t^2}}\right)^2\\\\ &=\left(\frac{1}{1+t^2}-\frac{t^2}{1+t^2}\right)^2+\left(\frac{2t}{1+t^2}\right)^2\\\\ &=\left(\frac{1-t^2}{1+t^2}\right)^2+\left(\frac{2t}{1+t^2}\right)^2\\\\ 1&=x^2+y^2 \end{align*}\] Look familiar?